### v1

# Determinant and Pfaffian formulas for particle annihilation

Abstract.
When particles on a line collide, they may annihilate—both are destroyed. Computing exact annihilation probabilities has been difficult because collisions reduce the particle count, while determinantal methods require a fixed count throughout. The ghost particle method, introduced in a companion paper for coalescence, keeps destroyed particles walking as invisible ghosts that restore the missing dimension. We apply this method to annihilation: when two particles annihilate, both trajectories continue as invisible walkers, yielding an exact determinantal formula that specifies the number of annihilations, where survivors end up, and where ghosts end up. For complete annihilation (no survivors), the formula simplifies to a Pfaffian—an algebraic relative of the determinant built from pairwise quantities—connecting to Pfaffian point process theory. The annihilation formula also yields results about coalescence: pairwise coalescence can be reinterpreted as complete annihilation, producing a Pfaffian coalescence formula. These formulas are exact for any finite initial configuration and apply to discrete lattice paths, birth-death chains, and continuous diffusions including Brownian motion.
Key words and phrases:
annihilating random walks, ghost particles, Lindström–Gessel–Viennot lemma, Pfaffian, determinant, lattice paths, interacting particle systems
2020 Mathematics Subject Classification:
Primary 05A15; Secondary 05A19, 15A15, 60C05, 60J65, 82C22
To Marek Bożejko,
who taught me that in creation and annihilation,
order is everything.
1. Introduction
1.1. The problem
Consider $n$ particles performing independent random walks on a line. When two particles meet, they annihilate: both are destroyed. This $A+A\to\emptyset$ reaction-diffusion model appears throughout statistical physics, modeling recombination kinetics, domain wall dynamics [GPTZ18], and pair annihilation [AH00]. What is the probability that exactly $k$ annihilations occur, with survivors reaching specified positions?
For non-colliding particles, exact probabilities are classical. The Karlin–McGregor theorem [KM59] and its combinatorial cousin, the Lindström–Gessel–Viennot (LGV) lemma [Lin73, GV85], express the probability that $n$ particles starting at positions $x_{1}<\cdots<x_{n}$ reach positions $y_{1},\ldots,y_{n}$ without colliding as a determinant:
$\Pr(\text{particles reach }y_{1},\ldots,y_{n}\text{ without colliding})=\det\bigl(p(x_{i}\to y_{j})\bigr)_{1\leq i,j\leq n},$
where $p(x_{i}\to y_{j})$ is the transition probability from $x_{i}$ to $y_{j}$.
When particles annihilate, the count decreases. After $k$ collisions, only n2kn-2k particles remain. The number of rows (initial particles) exceeds the number of columns (final particles), and there is no square matrix to write down. This dimensional mismatch places annihilating systems outside the scope of the classical LGV framework.
1.2. The ghost method
The companion paper [Śni26a] introduced ghost particles for coalescence: when two particles merge, one heir and one ghost emerge, keeping the entity count at $n$ and enabling determinantal formulas.
For annihilation, the same idea applies with a twist: when two particles annihilate, an ordered pair of ghosts emerges from the collision point—invisible walkers that perform independent random walks onward, without interacting with anything. The ghosts are anonymous: only the twin pairing is preserved (Sections˜2.3 and 1.6). The entity count—survivors plus ghosts—remains exactly $n$, restoring the dimensional structure needed for determinantal methods. LABEL:fig:intro-annihilation illustrates: particles $2$ and $3$ annihilate at $c$, and a ghost pair emerges, while particles $1$ and $4$ survive.
1.3. Main results
1.3.1. The annihilation formula
Consider $n$ particles starting at $x_{1}<\cdots<x_{n}$. Suppose $k$ collisions occur, producing s=n2ks=n-2k survivors at positions $y_{1}<\cdots<y_{s}$ and $k$ ghost pairs. The ghost pairs are physically indistinguishable—they do not remember which particles created them—so we label them $1,\ldots,k$ by sampling a uniform random numbering. Ghost pair $j$ then has positions $(a_{j},b_{j})$.
Theorem 1.1 (Annihilation formula with ghosts).
For a fixed final state—survivor positions $y_{1},\ldots,y_{s}$ and ghost pair positions $(a_{1},b_{1}),\ldots,(a_{k},b_{k})$—the probability is:
$\Pr=\frac{1}{k!}\Bigl[\prod_{j=1}^{k}t_{j}^{\varepsilon_{j}}\Bigr]\det(M),$
where $\varepsilon_{j}=+1$ if $a_{j}\leq b_{j}$ (ghost pair ordered left-right), 1-1 if $a_{j}>b_{j}$ (right-left).
The notation $\prod_{j}t_{j}^{\varepsilon_{j}}]$ means: expand the determinant in formal variables $t_{j}^{+}$ and $t_{j}^{-}$, then extract the coefficient where each ghost pair $j$ contributes $t_{j}^{+}$ if $\varepsilon_{j}=+1$, or $t_{j}^{-}$ if $\varepsilon_{j}=-1$. The factor $1/k!$ is the probability of the specific numbering $1,\ldots,k$.
The $n\times n$ matrix $M$ has rows indexed by initial particles and columns indexed by final entities (survivors and ghosts). Survivor columns contain transition probabilities $p(x_{i}\to y_{\ell})$. Ghost columns additionally carry formal variables.
1.3.2. Example: four particles, one collision
Four particles start at $x_{1}<x_{2}<x_{3}<x_{4}$, with one collision ($k=1$). Two survivors end at $y_{1}<y_{2}$ and one ghost pair at positions $a$ and bb—as in LABEL:fig:intro-annihilation.
The $4\times 4$ matrix $M$ has two survivor columns and two ghost columns (one pair). Since $k=1$, we drop the pair index, writing $\tau^{\pm}_{I}$ for $\tau^{(1)\pm}_{I}$ and $t^{\pm}$ for $t_{1}^{\pm}$:
$M=\begin{pmatrix}p(x_{1}\to y_{1})&p(x_{1}\to y_{2})&\tau^{+}_{1}p(x_{1}\to a)&\tau^{-}_{1}p(x_{1}\to b)\\ p(x_{2}\to y_{1})&p(x_{2}\to y_{2})&\tau^{+}_{2}p(x_{2}\to a)&\tau^{-}_{2}p(x_{2}\to b)\\ p(x_{3}\to y_{1})&p(x_{3}\to y_{2})&\tau^{+}_{3}p(x_{3}\to a)&\tau^{-}_{3}p(x_{3}\to b)\\ p(x_{4}\to y_{1})&p(x_{4}\to y_{2})&\tau^{+}_{4}p(x_{4}\to a)&\tau^{-}_{4}p(x_{4}\to b)\end{pmatrix}$.
The first two columns are plain transition probabilities (survivor columns). The last two columns carry additional formal variables $\tau^{\pm}_{I}$ indexed by the row. These variables obey the multiplication rule:
$\tau^{+}_{I}\,\tau^{-}_{J}=\begin{cases}-t^{+}&\text{if }I>J,\\ t^{-}&\text{if }I<J.\end{cases}$
To extract the annihilation probability, expand $\det(M)$ by Laplace expansion along the two ghost columns. Each term assigns a pair of particles $\{I,J\}$ with $I<J$ to the ghost pair and the remaining two to the survivors. Working out the product rules and extracting $t^{\varepsilon}]$ for either sign gives a single formula:
(1.1) $\Pr=\frac{1}{1!}\sum_{\begin{subarray}{c}I<J\\ \{I,J\}\subset\{1,2,3,4\}\end{subarray}}(-1)^{I+J+1}\,p\bigl(x_{I}\to\max(a,b)\bigr)\,p\bigl(x_{J}\to\min(a,b)\bigr)\\ \cdot\det\bigl(p(x_{i}\to y_{\ell})\bigr)_{\begin{subarray}{c}i\in\{1,\ldots,4\}\setminus\{I,J\}\\ \ell\in\{1,2\}\end{subarray}},$
where the sign $(-1)^{I+J+1}$ comes from the Laplace expansion along columns $3$ and $4$, and each $2\times 2$ determinant is a Karlin–McGregor probability for the two surviving particles. The spatial ordering is always reversed: the originally rightward particle $J$ ends at $\min(a,b)$ and the originally leftward particle $I$ at $\max(a,b)$.
1.3.3. The Pfaffian connection
For complete annihilation (all particles destroyed), the formula simplifies dramatically. When $n=2k$ particles annihilate completely, the probability is a Pfaffian of pairwise quantities:
$P=\operatorname{Pf}(A),$
where $A_{IJ}$ is the pairwise annihilation probability for particles $I$ and $J$. This connects to the Pfaffian point process theory for annihilating and coalescing particles [TZ11, GPTZ18]; Section˜5 derives the Pfaffian structure via a purely combinatorial route.
1.3.4. Relation to the LGV lemma
The annihilation formula generalizes the Karlin–McGregor / Lindström–Gessel–Viennot (LGV) determinant, just as the coalescence formula does [Śni26a]. In both cases, the proof uses a sign-reversing involution that swaps path segments at the first wrong crossing—but in annihilation, “wrong” means “crossing that violates the prescribed collision pattern” rather than any crossing at all. When no collisions occur (all particles survive, no ghosts), the formula reduces to the standard LGV determinant. See [Śni26a] for the full discussion of the LGV connection, including the relationship to Stembridge’s $D$-compatibility [Ste90] and Fisher’s vicious walkers [Fis84].
1.4. Prior work
Exact distributional results for annihilating particles have been obtained through two main approaches. Tribe and Zaboronski [TZ11] proved Pfaffian point process structure for annihilating Brownian motions under the maximal entrance law: their Pfaffian structure (Theorem 2 in [TZ11]) is exact, while their $n$-point density formula (Theorem 1) is asymptotic for large $t$. Garrod, Poplavskyi, Tribe, and Zaboronski [GPTZ18] extended the Pfaffian structure to discrete time and deterministic initial conditions. The interparticle distribution function (IPDF) method, developed by Doering and ben-Avraham [DA88] for diffusion-limited coalescence and extended to full correlation hierarchies in [Avr98], also applies to annihilation via the method of intervals [AH00]. Each of these approaches relies on analytic structure specific to the process at hand.
Our approach is combinatorial, as in the companion paper [Śni26a]: ghosts maintain a fixed entity count, enabling LGV-style determinants that apply uniformly to lattice paths, birth-death chains, and Brownian motion. Instead of the Pfaffian structure arising from infinite initial data, we have determinantal formulas that are exact for finite initial configurations. In particular, the ghost formula gives exact probabilities specifying the number of annihilations, where survivors end up, and where ghosts end up—going beyond the aggregate statistics (particle density, gap distributions) that analytic methods typically provide.
1.5. Applications
1.5.1. Domain wall dynamics
In the one-dimensional Ising–Glauber model, a spin chain evolves by single-spin flips at rate proportional to the energy change. Domain walls—boundaries between spin-up and spin-down regions—perform random walks on the dual lattice [GPTZ18]. When two domain walls meet, the intervening region disappears and both walls are destroyed: this is precisely the $A+A\to\emptyset$ annihilation dynamics.
The annihilation formula gives exact probabilities for the evolution of domain wall configurations. Given $n$ domain walls at positions $x_{1}<\cdots<x_{n}$, the formula computes the probability that $k$ pairs annihilate, with survivors reaching positions $y_{1}<\cdots<y_{n-2k}$ and ghost pairs (recording annihilation locations) at specified positions. This goes beyond the density and correlation functions typically computed for Ising–Glauber dynamics.
1.5.2. Reaction-diffusion systems
The $A+A\to\emptyset$ reaction models diffusion-limited pair annihilation: particles diffuse independently until two meet and react, destroying both. This appears in recombination kinetics (electron-hole pairs in semiconductors), chemical reactions at low concentrations, and population dynamics with fatal encounters [AH00].
The ghost formula provides exact finite-time probabilities for specific outcomes: not just how many particles survive, but which positions survivors and ghost pairs reach.
1.6. Scope and limitations
The introduction presents formulas for discrete random walks, but the method applies more broadly:
  • Continuous processes. The discrete framework (Section˜2 develops it for arbitrary planar directed acyclic graphs) extends to Brownian motion, birth-death chains, and other processes satisfying the Karlin–McGregor assumptions [KM59] (strong Markov property, meeting times are stopping times). The full continuous-time treatment is developed in the companion paper [Śni26a].
  • Ghost anonymity. Our formula treats ghost pairs as anonymous: they do not remember which initial particles produced them. Can this be refined to prescribe which particles annihilate—that is, can ghosts remember their origin? Appendix˜A provides computational evidence that no such refinement exists: the prescribed annihilation probability cannot be expressed as any linear combination of Karlin–McGregor products, even allowing arbitrary rational coefficients. Ghost anonymity is not merely a bookkeeping convenience but appears essential for the formula to exist.
1.7. Companion papers and organization
The ghost method applies equally to coalescence ($A+A\to A$). The companion paper [Śni26a] develops the coalescence formula, a ghost-free coalescence determinant (integrating out ghost positions), and the full continuous-time treatment. The two papers share the same proof architecture—sign-reversing involution via segment swap—but differ in the ghost structure: annihilation produces ghost pairs with paired formal variables, while coalescence produces single ghosts with a staircase pattern. The Pfaffian pairwise coalescence formula (Section˜5) bridges the two models: it is proved by converting pairwise coalescence to complete annihilation via parity reclassification.
Another companion paper [Śni26] develops the ghost framework into concrete gap statistics for coalescing particles: the Rayleigh distribution for the single-gap law, an explicit joint distribution for consecutive gaps (revealing negative correlation: a large gap makes adjacent gaps smaller), a generalization of Warren’s determinantal formula [War07] from Brownian motion to all skip-free processes, and a Pfaffian empty-interval formula that recovers known results of ben-Avraham [Avr98] from the ghost framework. Voter model duality bridges the ghost framework to the Pfaffian point process theory of Tribe and Zaboronski [TZ11] and Garrod, Poplavskyi, Tribe, and Zaboronski [GPTZ18].
The paper is organized as follows. Section˜2 sets up the spacetime graph framework and formalizes the annihilation model (collision diagrams, ghost pairs, performances). Section˜3 defines the matrix $M$ with its formal variables and states the annihilation formula. The proof occupies Section˜4: castings assign particles to roles, attribution and rehearsal connect castings to performances, and a sign-reversing involution cancels all non-physical contributions. Finally, Section˜5 converts pairwise coalescence to complete annihilation, deriving a Pfaffian formula that connects to Pfaffian point process theory [TZ11, GPTZ18].
2. Setup
The annihilation formula holds for random walks on $\mathbb{Z}$, Brownian motion on $\mathbb{R}$, and birth-death chains on arbitrary state spaces. Following the companion paper [Śni26a], we work with spacetime graphs—a combinatorial abstraction that captures two structural properties common to all these settings: planarity (paths with swapped endpoints must cross, and non-adjacent particles cannot meet without an intermediate particle involved) and weight-preserving segment swap (exchanging path segments at crossings preserves total weight). For discrete models, the spacetime graph is literal; for continuous processes, the combinatorial structure is identical but the proof uses measure-theoretic arguments [Śni26a].
2.1. Spacetime graphs
Definition 2.1 (Spacetime graph).
A spacetime graph is a directed acyclic graph $D=(V,E)$ with edge weights $w\colon E\to R$, where $R$ is a commutative ring. The DAG structure induces a time ordering: $u\lessdot v$ if there is a directed path from $u$ to $v$.
Definition 2.2 (Paths and weights).
A path from $x$ to $y$ is a sequence of vertices $(v_{0},\ldots,v_{\ell})$ with $v_{0}=x$, $v_{\ell}=y$, and each $(v_{i},v_{i+1})\in E$. The weight of a path is $w(P)=\prod_{i}w(v_{i}\to v_{i+1})$. The path generating function is
$W(x\to y)=\sum_{P:x\to y}w(P).$
2.2. Planarity
Definition 2.3 (Source and target sets).
The source set $\mathcal{X}\subseteq V$ and target set $\mathcal{Y}\subseteq V$ are each equipped with a linear order $\prec$.
Definition 2.4 (Planar configuration).
The pair $(\mathcal{X},\mathcal{Y})$ is planar if:
  1. (P1)
    Crossing property. For $x\prec x^{\prime}$ in $\mathcal{X}$ and $y^{\prime}\prec y$ in $\mathcal{Y}$ (targets swapped), every path from $x$ to $y$ intersects every path from $x^{\prime}$ to $y^{\prime}$.
  2. (P2)
    Consecutive collision property. For $x\prec x^{\prime}\prec x^{\prime\prime}$ in $\mathcal{X}$, if paths from $x$ and $x^{\prime\prime}$ meet at vertex $v$, then every path from $x^{\prime}$ must pass through $v$ or intersect one of those paths before $v$.
The crossing property (P1) is Stembridge’s $D$-compatibility [Ste90]: paths with swapped endpoints must meet. The consecutive collision property (P2), introduced in the companion paper [Śni26a], ensures that non-adjacent particles cannot collide without involving intermediate ones. Both properties hold for lattice paths, random walks on $\mathbb{Z}$, birth-death chains, and Brownian motion; see [Śni26a] for verification. The consecutive collision property is essential for the sign-reversing involution: it ensures that when two paths cross, the corresponding particles are adjacent in the active set (Section˜4.6).
2.3. The annihilation model
Fix $n$ source vertices $x_{1}\prec x_{2}\prec\cdots\prec x_{n}$ in $\mathcal{X}$ and a set of target vertices in $\mathcal{Y}$. Each source vertex represents an initial particle; a path from $x_{I}$ to a target vertex $y$ represents the particle’s trajectory, with weight $W(x_{I}\to y)$ equal to the path generating function on the spacetime graph $D$.
When two particles occupy the same vertex, they annihilate: both are destroyed and a twin pair of ghosts emerges from the collision point. The ghosts do not remember which particles produced them; only the twin pairing is preserved as the ghosts drift apart.
With $k$ collisions, we have s=n2ks=n-2k survivors and $2k$ ghosts (forming $k$ pairs), preserving the total count of $n$ entities. The rest of this section makes this precise: Section˜2.4 defines actors and roles, Section˜2.5 formalizes collision diagrams, and Section˜2.6 assembles these into performances.
2.4. Actors and roles
Definition 2.5 (Actors).
The actor set $\mathcal{A}=\{1,\ldots,n\}$ indexes the initial particles. Actor $I\in\mathcal{A}$ starts at position $x_{I}$.
Definition 2.6 (Roles).
The role set (final entities) is:
$\mathcal{R}=\mathcal{S}\cup\mathcal{G},$
where:
  • $\mathcal{S}=\{1,\ldots,s\}$: survivor slots (with s=n2ks=n-2k);
  • $\mathcal{G}=\{1,\ldots,k\}\times\{1,2\}$: ghost slots.
The cardinality is $s+2k=n$, matching $|\mathcal{A}|$.
Each role $f\in\mathcal{R}$ has a final position $y_{f}$:
  • Survivor slot $\ell\in\{1,\ldots,s\}$: position $y_{\ell}$ (with $y_{1}\prec\cdots\prec y_{s}$);
  • Ghost slot $(j,m)\in\{1,\ldots,k\}\times\{1,2\}$: position $y_{(j,m)}$.
We write $a_{j}=y_{(j,1)}$ and $b_{j}=y_{(j,2)}$ as shorthand for the positions of the two ghosts in pair $j$.
2.5. Collision diagrams
Definition 2.7 (Collision diagram).
A collision diagram is a graph embedded in the spacetime graph $D$ with three types of vertices:
  • Initial vertices $x_{1},\ldots,x_{n}$: each has in-degree 0 and out-degree $1$;
  • Collision vertices: each has out-degree 0 or $1$ and even total degree. At a collision vertex with $m$ incoming edges and out-degree $d\in\{0,1\}$, the collision produces (md)/2(m-d)/2 ghost pairs (and one outgoing edge if $d=1$);
  • Survivor vertices $y_{1},\ldots,y_{s}$: each has in-degree $1$ and out-degree 0.
Each initial vertex has a directed path to either a collision vertex or a survivor vertex. Distinct paths meet only at collision vertices (no crossings between collisions). The collision diagram records which collisions occur and where, but not what happens after each collision; the ghost paths are specified separately as part of the performance (Definition˜2.8).
2.6. Performances
Definition 2.8 (Performance).
An annihilation performance $\mathcal{P}$ specifies:
  • A collision diagram: which particles annihilated and where;
  • Final positions for each survivor;
  • At each collision vertex $v$: an ordered list of ghost pairs produced at $v$. A vertex with $m$ incoming edges and outdegree $d$ produces (md)/2(m-d)/2 ghost pairs;
  • A global numbering $j=1,\ldots,k$ of all ghost pairs across all vertices;
  • For each ghost pair $j$: an ordered pair of ghost paths from the collision vertex to positions $a_{j}$ and $b_{j}$.
The weight of a performance is
[…]
where $\Gamma_{j,1}$ is the ghost path to position $a_{j}=y_{(j,1)}$ and $\Gamma_{j,2}$ is the ghost path to position $b_{j}=y_{(j,2)}$.
Performances are role-based: they specify what happened (collisions, final positions) without tracking which initial particle ended where. Ghost pairs are physically indistinguishable—they do not remember which particles created them. Mathematics does not handle unordered collections as naturally as ordered ones (for instance, a probabilist wants the final state to be a random variable taking values in a space of ordered tuples). We therefore convert the collection of ghost pairs into an ordered sequence by sampling a uniform random numbering. The factor $1/k!$ is the probability of any specific numbering; summing over all $k!$ numberings of a fixed physical outcome recovers its probability.
The global numbering determines the final state: ghost pair $j$ occupies slots $(j,1)$ and $(j,2)$ with positions $a_{j}$ and $b_{j}$. The proof (Section˜4) introduces the complementary actor-based perspective via castings.
LABEL:fig:annihilation-performance shows a larger example with five particles and two collisions on the $\mathbb{Z}^{2}$ lattice.
2.7. Final state
Definition 2.9 (Ghost sign).
For ghost pair $j$, the ghost sign is:
$\varepsilon_{j}=\begin{cases}+1&\text{if }a_{j}\preceq b_{j},\\ -1&\text{if }b_{j}\prec a_{j}.\end{cases}$
When $a_{j}=b_{j}$ (ghosts at the same final position), we have $\varepsilon_{j}=+1$ by convention. For continuous state spaces, this case has probability zero.
Definition 2.10 (Final state).
The final state $\mathcal{F}$ specifies:
  • The number $k$ of collisions (with $2k\leq n$);
  • Final positions $y_{f}$ for each $f\in\mathcal{R}$.
The sign of the final state is the product of ghost signs: $\operatorname{sgn}\mathcal{F}=\prod_{j=1}^{k}\varepsilon_{j}$.
3. The annihilation formula
3.1. Formal variables
The determinant $\det(M)$ sums over all bijections $\pi\colon\mathcal{A}\to\mathcal{R}$—all possible assignments of initial particles to final entities. Only those assignments consistent with the ghost configuration should contribute. We introduce formal variables that filter the Leibniz expansion, selecting the correct terms via coefficient extraction.
For each ghost pair $j\in\{1,\ldots,k\}$, introduce a formal parameter $t_{j}$ and formal variables $\tau^{(j)+}_{I}$, $\tau^{(j)-}_{I}$ for each particle $I$, subject to the relations (for $I\neq J$):
(3.1) $\tau^{(j)+}_{I}\,\tau^{(j)-}_{J}=\begin{cases}-t_{j}^{+}&\text{if }I>J,\\ t_{j}^{-}&\text{if }I<J.\end{cases}$
(The case $I=J$ does not arise: in the Leibniz expansion of $\det(M)$, each bijection assigns distinct particles to the two ghost slots.)
Concretely, the product rule (3.1) assigns $t_{j}^{+}$ when the higher-indexed particle goes to slot $(j,1)$ and $t_{j}^{-}$ when the lower-indexed one does, so extracting $t_{j}^{+}]$ selects exactly those terms where the spatial ordering is reversed.
3.2. The matrix
Define the $n\times n$ matrix $M$ with:
  • Rows indexed by particles $I\in\mathcal{A}$;
  • Columns indexed by roles $f\in\mathcal{R}$: survivor slots $1,\ldots,s$ followed by ghost slots $(1,1),(1,2),\ldots,(k,1),(k,2)$.
Entries:
[…]
3.3. The theorem
Theorem 3.1 (Annihilation formula with ghosts).
For a fixed final state $\mathcal{F}$, the total weight of performances is:
$Z=\frac{1}{k!}\Bigl[\prod_{j=1}^{k}t_{j}^{\varepsilon_{j}}\Bigr]\det(M).$
When $k=0$ (no collisions), the formula reduces to the Karlin–McGregor determinant $Z=\det(M)$.
Here $\prod_{j}t_{j}^{\varepsilon_{j}}]$ denotes coefficient extraction: for each ghost pair $j$, extract the coefficient of $t_{j}^{+}$ if $\varepsilon_{j}=+1$ (slot $(j,1)$ has smaller or equal position), or $t_{j}^{-}$ if $\varepsilon_{j}=-1$ (slot $(j,2)$ has smaller position).
4. Proof of the annihilation formula
This proof parallels the coalescence proof in [Śni26a]: identical four-part structure (castings, attribution, rehearsal, sign-reversing involution) and identical involution mechanism (segment swap at the first wrong crossing). The annihilation proof is simpler because annihilation destroys both particles at a collision—no heir continues—reducing the case analysis in the involution. Nevertheless, ghost pairs differ enough from single ghosts that the success criterion, attribution map, and involution case analysis all require reworking. The present proof is self-contained; familiarity with the companion paper is not required.
4.1. Performances vs. castings
Throughout the proof, we fix a final state $\mathcal{F}$ (Section˜2.7): the number $k$ of collisions, survivor positions $y_{1},\ldots,y_{s}$, and ghost pair positions $(a_{j},b_{j})$ for $j=1,\ldots,k$. The generating function $Z=Z_{\mathcal{F}}$ counts performances for $\mathcal{F}$: complete descriptions of what collisions occur and where the final entities end up.
4.1.1. Performances
A performance is role-based—it specifies outcomes without tracking which particle “becomes” which final entity. In probabilistic terms, each performance is an elementary event in the sample space.
Consider LABEL:fig:intro-annihilation: four particles start at $x_{1}<x_{2}<x_{3}<x_{4}$, particles $2$ and $3$ collide at $c$, and the system ends with two survivors at $y_{1}<y_{2}$ and a ghost pair at $(a,b)$. The role-based description records:
  • A collision occurred at spacetime point $c$.
  • The ghost pair emerged from that collision, reaching positions $a$ and $b$.
  • Two particles survived, reaching $y_{1}$ and $y_{2}$.
  • The paths taken by each entity.
This description says nothing about which of particles $2$ and $3$ ended at which ghost position—only that the collision produced a ghost pair. In theater terms: the script specifies where each character should stand at the final scene, but does not specify which actor plays which character.
4.1.2. The determinant perspective
The determinant $\det(M)$ gives something different: a signed sum over castings—assignments of particles to final positions via non-interacting paths. A casting is actor-based: it tracks each particle’s complete trajectory. LABEL:fig:successful-casting shows a successful casting where line styles persist through the collision, revealing who went where.
The proof establishes a bijection between performances and a subset of castings (the “successful” ones), while the remaining “failed” castings cancel in pairs (LABEL:fig:failed-casting).
4.1.3. Outline of the argument
The argument has four parts:
  1. (1)
    Castings from the determinant. The Leibniz expansion produces castings; coefficient extraction selects candidate castings consistent with the ghost configuration.
  2. (2)
    Attribution. Every performance determines a casting by tracking which particle ends where. Attribution always produces a candidate casting.
  3. (3)
    Rehearsal. We attempt to interpret each casting as a performance by scanning crossings in temporal order. This may succeed (“successful casting”) or fail (“failed casting”).
  4. (4)
    The sign-reversing involution. Failed castings pair up via segment swap and cancel. Only successful castings remain.
4.2. Castings
The Leibniz formula expands the determinant as a sum over bijections $\pi\colon\mathcal{A}\to\mathcal{R}$:
$\det(M)=\sum_{\pi}\operatorname{sgn}(\pi)\prod_{I\in\mathcal{A}}M_{I,\pi(I)}.$
Each matrix entry $M_{I,f}$ is either a path generating function $W(x_{I}\to y_{f})$ (for survivor columns) or $\tau^{(j)\pm}_{I}W(x_{I}\to y_{(j,m)})$ (for ghost columns).
Definition 4.1 (Casting).
A casting $(\pi,\mathbf{P})$ consists of:
  • A bijection $\pi\colon\mathcal{A}\to\mathcal{R}$;
  • A path family $\mathbf{P}=\{P_{I}\}_{I\in\mathcal{A}}$ where $P_{I}$ goes from $x_{I}$ to $y_{\pi(I)}$.
The weight of a casting is
$w(\mathcal{C})=\frac{1}{k!}\prod_{I\in\mathcal{A}}w(P_{I}).$
Crucially, castings are pure geometry: the paths are non-interacting. They may cross freely—a crossing is simply a shared vertex with no physical consequence.
Definition 4.2 (Candidate bijection).
A bijection $\pi\colon\mathcal{A}\to\mathcal{R}$ is a candidate for $\mathcal{F}$ if, for each ghost pair $j\in\{1,\ldots,k\}$:
  • when $\varepsilon_{j}=+1$: $\pi^{-1}((j,1))>\pi^{-1}((j,2))$;
  • when $\varepsilon_{j}=-1$: $\pi^{-1}((j,1))<\pi^{-1}((j,2))$.
In other words, when $a_{j}\preceq b_{j}$ (ghost pair ordered left-to-right), the higher-indexed particle maps to the leftward position $a_{j}$.
Candidacy arises from the coefficient extraction. Each ghost pair $j$ contributes two columns to $\det(M)$: column $(j,1)$ with entries $\tau^{(j)+}_{I}$ and column $(j,2)$ with entries $\tau^{(j)-}_{I}$. When a bijection $\pi$ sends particles $I$ to $(j,1)$ and $J$ to $(j,2)$, the product rule (3.1) yields $\tau^{(j)+}_{I}\tau^{(j)-}_{J}=-t_{j}^{+}$ if $I>J$ and $t_{j}^{-}$ if $I<J$. Extracting $t_{j}^{\varepsilon_{j}}$ selects those bijections satisfying the candidacy condition.
Write $\Pi_{\mathcal{F}}$ for the set of candidate bijections. After coefficient extraction, only candidates survive. Each candidate $\pi$ contributes a sign from the formal variable product (determined by the ordering of actors within each ghost pair) in addition to $\operatorname{sgn}(\pi)$:
$\Bigl[\prod_{j}t_{j}^{\varepsilon_{j}}\Bigr]\det(M)=\sum_{\pi\in\Pi_{\mathcal{F}}}(\text{formal sign})\cdot\operatorname{sgn}(\pi)\prod_{I}W(x_{I}\to y_{\pi(I)}).$
The sign analysis in Section˜4.8 shows that these signs combine to give +1+1 for each successful casting.
4.3. Attribution
A performance specifies collision locations and final positions but not identities. Attribution constructs the underlying casting by tracking which particle ends where. We first describe what happens at a single collision, then explain how to combine these local operations.
4.3.1. The two-particle case
At a pairwise annihilation, two path segments arrive and two ghost paths depart. The swap principle determines the gluing:
Principle 4.3 (Swap principle).
At a collision between two particles: the particle that starts more to the left continues to the ghost that ultimately ends more to the right; the particle that starts more to the right continues to the ghost that ultimately ends more to the left.
LABEL:fig:simplified-annihilation illustrates the case $\varepsilon_{j}=+1$ (that is, $a_{j}\preceq b_{j}$). Two ghost paths depart the collision vertex toward positions $a_{j}$ and $b_{j}$, see LABEL:fig:simplified-annihilation-a. The swap principle glues each incoming path to an outgoing ghost path, see LABEL:fig:simplified-annihilation-b: particle $J$ (right, larger index) is glued to the leftward ghost at $a_{j}$, and particle $I$ (left, smaller index) is glued to the rightward ghost at $b_{j}$. The case $\varepsilon_{j}=-1$ (meaning $b_{j}\prec a_{j}$) is symmetric: the labels $a_{j}$ and $b_{j}$ swap sides, so particle $J$ again goes to the leftward position and particle $I$ to the rightward one. In both cases, the spatial ordering is reversed.
4.3.2. From local to global
Attribution takes a performance as input and produces a casting $(\pi,\mathbf{P})$. For each initial particle $I$, follow its path to the collision vertex (if any) and apply the swap principle to determine which ghost path it joins. The bijection $\pi$ combines local and global structure: the second coordinate of $\pi(I)=(j,m)$ comes from the swap principle (which assigns $I$ to either the first or second path in the ordered pair, giving $m\in\{1,2\}$), while the first coordinate $j$ comes from the performance’s global numbering of ghost pairs.
Definition 4.4 (Attribution).
Attribution constructs a casting $(\pi,\mathbf{P})$ from a performance by applying the swap principle at each collision vertex. For each initial particle $I$:
  • If $I$ collides, the swap principle glues its incoming path to an outgoing ghost path, determining the endpoint $\pi(I)\in\mathcal{G}$ and the glued path $P_{I}$.
  • If $I$ survives, $\pi(I)\in\mathcal{S}$ and $P_{I}$ is the path from $x_{I}$ to the survivor position (this path avoids all collision vertices).
The output is a casting $(\pi,(P_{I})_{I\in\mathcal{A}})$.
Since both colliding particles are destroyed, each particle participates in at most one collision. Attribution therefore applies the swap principle independently at each collision vertex—no chaining arises (unlike coalescence, where the heir continues and may collide again).
At a high-indegree vertex where $m>2$ particles $I_{1}<I_{2}<\cdots<I_{m}$ meet simultaneously, we glue the first pair $(I_{1},I_{2})$ to the first ordered pair of ghosts in the vertex’s list, the second pair $(I_{3},I_{4})$ to the second ordered pair, and so on. If $m$ is odd, the last incoming particle $I_{m}$ is glued to the outgoing edge. Each pair gluing applies the swap principle. (Section˜4.6, page 4.6, gives the detailed treatment.)
4.3.3. Example
Example 4.5 ($n=4$, $k=1$).
In LABEL:fig:intro-annihilation, particles $2$ and 33 ($I=2<3=J$) collide at $c$. The ghost pair has $a\preceq b$, so $\varepsilon=+1$. The swap principle glues particle 33 ($J$, right) to the leftward ghost at $a$, and particle 22 (II, left) to the rightward ghost at $b$. Survivors $1$ and $4$ pass through the collision diagram to their final positions. The resulting casting has $\pi(1)=1$ (survivor slot at $y_{1}$), $\pi(2)=(1,2)$, $\pi(3)=(1,1)$, $\pi(4)=2$ (survivor slot at $y_{2}$), with glued paths:
  • $P_{1}$: from $x_{1}$ to $y_{1}$ (survivor);
  • $P_{2}$: from $x_{2}$ to $c$, then to $b$ (ghost);
  • $P_{3}$: from $x_{3}$ to $c$, then to $a$ (ghost);
  • $P_{4}$: from $x_{4}$ to $y_{2}$ (survivor).
This is the successful casting shown in LABEL:fig:successful-casting.
4.3.4. Candidacy emerges
Proposition 4.6.
Let $(\pi,\mathbf{P})$ be the casting from attributing a performance. Then $\pi$ is a candidate.
Proof.
At the collision creating ghost pair $j$, let particles $I<J$ collide. When $a_{j}\preceq b_{j}$ (so $\varepsilon_{j}=+1$), the swap principle assigns $\pi(J)=(j,1)$ (position $a_{j}$, the leftward ghost) and $\pi(I)=(j,2)$ (position $b_{j}$, rightward). Then $\pi^{-1}((j,1))=J>I=\pi^{-1}((j,2))$, matching the candidacy condition for $\varepsilon_{j}=+1$. When $b_{j}\prec a_{j}$ (so $\varepsilon_{j}=-1$), the swap principle assigns $\pi(I)=(j,1)$ (position $a_{j}$, now rightward) and $\pi(J)=(j,2)$ (position $b_{j}$, leftward). Then $\pi^{-1}((j,1))=I<J=\pi^{-1}((j,2))$, matching the candidacy condition for $\varepsilon_{j}=-1$. ∎
4.4. Rehearsal
Rehearsal reverses attribution: given a candidate casting $(\pi,\mathbf{P})$, we attempt to interpret it as a performance by scanning crossings in temporal order.
The rehearsal algorithm:
  1. (1)
    Initialize: all particles are active.
  2. (2)
    Find the first crossing among active paths. If none exist, success.
  3. (3)
    Test whether this crossing is a valid annihilation (see below).
    • If yes: record the collision, deactivate both particles, continue from step 2.
    • If no: failure—the casting is “failed.”
Since both participants are deactivated at each valid collision, each particle is processed at most once.
4.4.1. Valid vs. spurious crossings
A crossing between particles $I<J$ is valid if both particles are destined for the same ghost pair: $\{\pi(I),\pi(J)\}=\{(j,1),(j,2)\}$ for some ghost pair jj—that is, the casting assigns $I$ and $J$ to the two slots of a single twin pair. If the constraint fails (one or both particles are destined for survivor slots, or they map to different ghost pairs), the crossing is spurious.
Definition 4.7 (Successful casting).
A candidate casting is successful if rehearsal processes all crossings as valid annihilations without encountering a spurious one.
LABEL:fig:castings illustrates both outcomes. In LABEL:sub@fig:successful-casting, particles $2$ and $3$ are assigned to the ghost pair: their crossing at $c$ is valid, and rehearsal succeeds. In LABEL:sub@fig:failed-casting, particles $1$ and $3$ are assigned to the ghost pair instead, but the first crossing involves particles $1$ and 22—a spurious crossing, since particle $2$ is destined for a survivor position.
Remark 4.8 (Dual interpretation).
The rehearsal algorithm has two equivalent readings:
  • As a bijection: does this casting correspond to a performance? Success produces a performance; failure means it does not.
  • As an involution: is this casting a fixed point of $\iota$? Success means fixed point; failure triggers segment swap at the first spurious crossing, pairing the casting with another failed casting (Section˜4.6).
The two viewpoints are unified: a casting is successful if and only if it is a fixed point of the global involution $\iota$.
Proposition 4.9.
Attribution and rehearsal are mutual inverses between performances and successful castings.
Proof.
Rehearsal $\circ$ Attribution $\operatorname{id}$. Let $\mathcal{P}$ be a performance. The casting $(\pi,\mathbf{P})=\mathfrak{A}(\mathcal{P})$ is a candidate (Section˜4.3.4). The crossings in $\mathbf{P}$ are exactly the collisions in $\mathcal{P}$, so rehearsal encounters them in the same temporal order and processes each as valid (the constraint $\{\pi(I),\pi(J)\}=\{(j,1),(j,2)\}$ holds by construction of attribution). Rehearsal therefore succeeds and reconstructs $\mathcal{P}$.
Attribution $\circ$ Rehearsal $\operatorname{id}$. Let $(\pi,\mathbf{P})$ be a successful casting and $\mathcal{P}$ the performance produced by rehearsal. We verify that $\mathfrak{A}(\mathcal{P})=(\pi,\mathbf{P})$ by checking each collision in temporal order.
At a collision between particles $I<J$ filling ghost pair $j$, candidacy ensures the swap principle assigns the same roles as $\pi$: if $\varepsilon_{j}=+1$, then $\pi^{-1}((j,1))>\pi^{-1}((j,2))$, so the rightward particle maps to $(j,1)$ and the leftward to (j,2)(j,2)—matching attribution. The case $\varepsilon_{j}=-1$ is symmetric. Since both particles deactivate, later collisions are unaffected. After all collisions, surviving particles retain their original paths and endpoints. Thus attribution recovers $(\pi,\mathbf{P})$. ∎
4.4.2. Crossings must exist
The rehearsal algorithm terminates either when it encounters a spurious crossing (failure) or when no crossings remain among active paths (success). We must verify that success can only occur when all ghost roles have been filled—otherwise the algorithm could “stall” with unfilled ghost roles.
Lemma 4.10 (Order preservation excludes ghosts).
If $\pi$ is a candidate bijection that preserves the spatial ordering ($I<J$ implies $y_{\pi(I)}\preceq y_{\pi(J)}$), then $\pi$ assigns every actor to a survivor slot.
Proof.
Suppose $\pi$ assigns actors to ghost pair $j$ with $\varepsilon_{j}=+1$ (meaning $a_{j}\preceq b_{j}$). Write $I=\pi^{-1}((j,1))$ and $J=\pi^{-1}((j,2))$. Candidacy requires $I>J$, but order preservation and $a_{j}\preceq b_{j}$ require $I<J$. Contradiction. The case $\varepsilon_{j}=-1$ is symmetric. ∎
Corollary 4.11 (Crossings exist).
If some active actor is assigned to a ghost slot in a candidate bijection, then at least two active paths must cross.
Proof.
If no active paths cross, the restriction of $\pi$ to active actors preserves spatial order. By Section˜4.4.2, no active actor maps to a ghost slot—contradiction. ∎
This ensures rehearsal cannot terminate with unfilled ghost roles: if any particle is still destined for a ghost position, crossings must exist among active paths.
Proposition 4.12.
Let $(\pi,\mathbf{P})$ be a successful candidate casting. The performance produced by rehearsal has the prescribed final state $\mathcal{F}$.
Proof.
By Section˜4.4.2, rehearsal continues as long as any active particle maps to a ghost slot. Each valid collision fills one ghost pair and removes both participants. After $k$ collisions, no active particle maps to a ghost slot; the remaining paths reach survivor positions without crossing. ∎
4.5. Segment swap
Failed castings cancel via segment swap.
Definition 4.13 (Segment swap).
Segment swap is a transformation on castings. Given a casting $(\pi,\mathbf{P})$ where paths $P_{I}$ and $P_{J}$ cross at vertex $v$, segment swap at $v$ produces a new casting $(\pi^{\prime},\mathbf{P}^{\prime})$:
  • $P_{I}^{\prime}$: prefix of $P_{I}$ from $x_{I}$ to $v$, then suffix of $P_{J}$ from $v$ to $y_{\pi(J)}$;
  • $P_{J}^{\prime}$: prefix of $P_{J}$ from $x_{J}$ to $v$, then suffix of $P_{I}$ from $v$ to $y_{\pi(I)}$.
The bijection updates to $\pi^{\prime}=(I\;J)\circ\pi$, swapping the destinations.
Lemma 4.14 (Swap properties).
Segment swap is:
  1. (i)
    an involution: swapping twice recovers the original;
  2. (ii)
    weight-preserving: $w(P_{I})\cdot w(P_{J})=w(P_{I}^{\prime})\cdot w(P_{J}^{\prime})$;
  3. (iii)
    sign-reversing: $\operatorname{sgn}(\pi^{\prime})=-\operatorname{sgn}(\pi)$.
Proof.
Property (i): after swapping at $v$, the paths $P_{I}^{\prime}$ and $P_{J}^{\prime}$ still cross at $v$. Swapping again restores the original suffixes.
Property (ii): each edge appears in exactly one path before and after the swap, so the product of weights is unchanged.
Property (iii): the bijection $\pi^{\prime}=(I\;J)\circ\pi$ differs from $\pi$ by a transposition, so $\operatorname{sgn}(\pi^{\prime})=-\operatorname{sgn}(\pi)$. ∎
LABEL:fig:involution-pair shows the involution in action on the failed casting from LABEL:fig:failed-casting. Segment swap at the spurious crossing exchanges the suffixes of particles $1$ and $2$, producing a paired failed casting with reversed permutation sign but identical weight. The crossing persists after the swap and remains spurious, confirming $\iota^{2}=\operatorname{id}$.
4.6. The sign-reversing involution
For each pair $(I,J)$ with $I<J$, the local involution $\iota_{I,J}$ acts on candidate castings. Given a candidate casting $\mathcal{C}=(\pi,\mathbf{P})$:
[…]
Lemma 4.15 (Fixed points are valid annihilations).
Let $(\pi,\mathbf{P})$ be a candidate casting whose paths $P_{I}$ and $P_{J}$ cross (with $I<J$ adjacent in the active set). The following are equivalent:
  • $(\pi,\mathbf{P})$ is a fixed point of $\iota_{I,J}$ (swapping at the first crossing would break candidacy);
  • the first crossing is a valid annihilation: $\{\pi(I),\pi(J)\}=\{(j,1),(j,2)\}$ for some ghost pair $j$.
Proof.
The swap exchanges the roles of $I$ and $J$: the new bijection $\pi^{\prime}$ satisfies $\pi^{\prime}(I)=\pi(J)$ and $\pi^{\prime}(J)=\pi(I)$.
Valid annihilation ($\{\pi(I),\pi(J)\}=\{(j,1),(j,2)\}$). Before swap, candidacy requires $\pi^{-1}((j,1))$ and $\pi^{-1}((j,2))$ to be ordered consistently with $\varepsilon_{j}$. After swap, their ordering is reversed. Candidacy is violated, so the casting is a fixed point.
Spurious crossing ($\{\pi(I),\pi(J)\}\neq\{(j,1),(j,2)\}$ for all $j$). We verify that $\pi^{\prime}$ is still a candidate. For each ghost pair $j$: if both slots $(j,1),(j,2)$ are outside $\{\pi(I),\pi(J)\}$, the swap does not affect pair $j$. Otherwise, exactly one slot belongs to $I$ or $J$, and the other belongs to some third particle $K$. Since $I$ and $J$ are adjacent, $K$ lies outside the interval $[I,J]$: either $K<I<J$ or $I<J<K$. In both cases, swapping $I$ and $J$ preserves the ordering of $K$ relative to the participant of pair $j$, so candidacy for pair $j$ is maintained. ∎
The consecutive collision property (P2) guarantees that the first crossing among active paths always involves adjacent particles, so the adjacency hypothesis is satisfied whenever the algorithm invokes this lemma.
Local involutions do not commute in general, but they do commute when applied to disjoint pairs of particles. Since annihilation deactivates both participants, any two pairs checked by the algorithm are necessarily disjoint.
Lemma 4.16 (Restricted commutativity).
If the algorithm checks both $(I,J)$ and $(K,L)$, then $\iota_{I,J}\circ\iota_{K,L}=\iota_{K,L}\circ\iota_{I,J}$.
Proof.
Any two checked pairs satisfy $\{K,L\}\cap\{I,J\}=\emptyset$. Disjoint local involutions act on different paths and commute. ∎
The global involution $\iota$ composes local involutions, processing crossings in temporal order. Successful castings are fixed points; failed castings pair up via segment swap.
Input: A candidate casting $\mathcal{C}=(\pi,\mathbf{P})$.
Output: A candidate casting—either $\mathcal{C}$ itself (fixed point) or a different casting $\mathcal{C}^{\prime}=\iota_{I,J}(\mathcal{C})$ with $\operatorname{sgn}(\pi^{\prime})=-\operatorname{sgn}(\pi)$.
Algorithm.
  1. (1)
    Initialize: all particles are active.
  2. (2)
    While active paths cross:
    1. (a)
      Find the pair $(I,J)$ with the earliest first crossing among active paths. (Tie-breaking: lexicographically minimal pair; this ensures restricted commutativity, Section˜4.6.)
    2. (b)
      Test: Is this a valid annihilation ($\{\pi(I),\pi(J)\}=\{(j,1),(j,2)\}$ for some $j$)?
    3. (c)
      If no (spurious crossing): return $\iota_{I,J}(\mathcal{C})$.
    4. (d)
      If yes (valid collision): deactivate both particles and continue.
  3. (3)
    Success—return $\mathcal{C}$ (a fixed point).
High-indegree vertices. When $m>2$ active particles meet at a single vertex, sort them by initial index: $I_{1}<I_{2}<\cdots<I_{m}$. Process consecutive pairs $(I_{1},I_{2})$, $(I_{3},I_{4})$, etc.; each pair produces an annihilation. If $m$ is odd, the last incoming particle $I_{m}$ is not paired at this vertex and remains active. This pairing rule is a convention: the only requirement is that attribution and rehearsal use the same rule, so that the bijection between successful castings and performances is preserved. Ghost anonymity ensures that the specific convention does not affect the annihilation weight.
Theorem 4.17 (Involution theorem).
The map $\iota$ is a weight-preserving, sign-reversing involution on candidate castings. Its fixed points are exactly the successful castings.
Proof.
We verify each claimed property.
Involution. Suppose $\iota(\mathcal{C})\neq\mathcal{C}$, meaning the algorithm swapped at pair $(I,J)$. Write $\mathcal{C}^{\prime}=\iota(\mathcal{C})$. Since all valid collisions before the swap point deactivated both participants, the active sets in $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are identical up to pair $(I,J)$. The swapped paths $P_{I}^{\prime}$, $P_{J}^{\prime}$ still cross at the same first vertex, and the constraint status (valid or spurious) depends on $\{\pi(I),\pi(J)\}$, which equals $\{\pi^{\prime}(J),\pi^{\prime}(I)\}$—the same unordered set. The crossing is still spurious, so $\iota$ swaps again: $\iota^{2}=\operatorname{id}$.
Weight preservation and sign reversal. These follow from Section˜4.5: the swap preserves total path weight and reverses permutation sign.
Fixed points. The algorithm is fixed precisely when every crossing encountered is a valid annihilation. By Section˜4.4.2, if any active particle maps to a ghost slot, crossings exist among active paths. By Section˜4.6, the order in which valid collisions are processed does not matter (any two checked pairs are disjoint, since annihilation deactivates both participants). The algorithm terminates when all ghost roles are filled and remaining paths do not cross: this is exactly a successful casting. ∎
4.7. The sign identity
Proposition 4.18 (Sign identity).
For any successful casting $(\pi,\mathbf{P})$:
$\operatorname{sgn}(\pi)=(-1)^{\#\{j:\varepsilon_{j}=+1\}}.$
Proof.
Each collision fills one ghost pair. Consider the collision filling ghost pair $j$: particles $I<J$ annihilate. By the consecutive collision property (P2), they are adjacent in the active set at the time of collision (any active particle $K$ with $I<K<J$ would cross one of their paths earlier, forcing a prior collision).
The roles $(j,1)$ and $(j,2)$ have canonical order $(j,1)<(j,2)$, and the actors have $I<J$. By the swap principle:
  • When $\varepsilon_{j}=+1$: $\pi(J)=(j,1)$ and $\pi(I)=(j,2)$. The smaller actor maps to the larger role and vice versa: one inversion, contributing 1-1.
  • When $\varepsilon_{j}=-1$: $\pi(I)=(j,1)$ and $\pi(J)=(j,2)$. Order is preserved: no inversion, contributing +1+1.
Survivors contribute no inversions: after all ghost pairs are filled, surviving actors map to survivor slots in spatial order (their paths do not cross). Since collisions remove disjoint adjacent pairs (with high-indegree vertices decomposed into consecutive binary annihilations as in Section˜4.6), contributions multiply independently. The total sign is $\operatorname{sgn}(\pi)=(-1)^{\#\{j:\varepsilon_{j}=+1\}}$. ∎
4.8. Completing the proof
Proof of Section˜3.3.
The coefficient extraction $\prod_{j}t_{j}^{\varepsilon_{j}}]$ selects candidate bijections. By Section˜4.6, the involution $\iota$ partitions candidate castings into:
  • Fixed points: successful castings;
  • Matched pairs: failed castings paired by segment swap.
Each matched pair consists of castings $(\pi,\mathbf{P})$ and $(\pi^{\prime},\mathbf{P}^{\prime})$ with equal path products $\prod_{I}w(P_{I})=\prod_{I}w(P_{I}^{\prime})$ (Section˜4.5(ii)) and $\operatorname{sgn}(\pi^{\prime})=-\operatorname{sgn}(\pi)$ (Section˜4.5(iii)). Since the $1/k!$ factor is the same for both, the casting weights are equal and their contributions to the determinant cancel.
Only successful castings survive. By Section˜4.4.1, these biject with performances. We verify that each contributes $1\cdot w(\mathcal{C})$ to $Z$:
Weight: For a successful casting $(\pi,\mathbf{P})$, the path product $\prod_{I}w(P_{I})$ runs over exactly the same paths as in the corresponding performance: diagram edges traversed by survivor paths, plus the ghost paths $\Gamma_{j,1}$, $\Gamma_{j,2}$ for each pair $j$. Since both casting and performance weights include the factor $1/k!$, the bijection is weight-preserving.
Sign cancellation: Each ghost pair $j$ contributes two sign factors. First, the Leibniz expansion contributes a factor to $\operatorname{sgn}(\pi)$; by Section˜4.7, this factor is 1-1 when $\varepsilon_{j}=+1$ and +1+1 when $\varepsilon_{j}=-1$. Second, the formal variable relations (Equation˜3.1) contribute a sign: candidacy with $\varepsilon_{j}=+1$ requires $\pi^{-1}((j,1))>\pi^{-1}((j,2))$, producing $t_{j}^{+}$ (a factor of 1-1); candidacy with $\varepsilon_{j}=-1$ requires $\pi^{-1}((j,1))<\pi^{-1}((j,2))$, producing $t_{j}^{-}$ (no extra sign). In both cases, the two factors are identical and their product is $(-1)^{2}=+1$ or $(+1)^{2}=+1$. Thus the net sign contribution from each ghost pair is +1+1, and the total contribution of each successful casting is $1\cdot w(\mathcal{C})$.
Conclusion: By Section˜4.4.1, successful castings and performances are in one-to-one correspondence. Since both have weight $\frac{1}{k!}\prod_{I}w(P_{I})$, the bijection preserves weight:
$\frac{1}{k!}\bigl[\textstyle\prod_{j}t_{j}^{\varepsilon_{j}}\bigr]\det(M)=\sum_{\text{successful}}w(\mathcal{C})=\sum_{\mathcal{P}}w(\mathcal{P})=Z.$
5. Pairwise coalescence and Pfaffians
This section demonstrates a surprising interplay between annihilation and coalescence: we prove that the total weight of pairwise coalescence is a Pfaffian of pairwise quantities, by converting the coalescence problem to a complete annihilation problem. The annihilation formula (Section˜3.3)—developed for $A+A\to\emptyset$—turns out to be the key tool for proving a result about $A+A\to A$. This connection arises from a parity reclassification that reinterprets coalescence events as annihilations.
5.1. The coalescence model
Before deriving the Pfaffian formula, we introduce the coalescence model from the companion paper [Śni26a]. We provide only the definitions necessary for the parity reclassification.
5.1.1. Coalescence dynamics
In the coalescence model ($A+A\to A$), when two particles collide one heir and one ghost emerge: the heir continues as a visible particle while the ghost becomes an invisible walker that no longer interacts with anything. This contrasts with annihilation ($A+A\to\emptyset$), where both entities become ghosts.
Definition 5.1 (Multiplicity).
Each entity carries a multiplicity: the number of original particles it represents. An initial particle has multiplicity $1$. At each coalescence, the heir has the sum of incoming multiplicities while the ghost has multiplicity 0.
Example 5.2.
If a particle with multiplicity $3$ (arising from earlier mergers) coalesces with a particle with multiplicity $2$, the heir has multiplicity $5$ and the ghost has multiplicity 0.
5.1.2. Coalescence performance
A coalescence performance specifies:
  • A genealogy forest: which particles merged and where;
  • Ghost paths: where each ghost traveled after emerging.
The weight of a performance is the product of all path weights. (Unlike annihilation, coalescence ghosts are individually identifiable—each emerges from a specific junction—so no random labeling factor is needed.) See [Śni26a] for the full treatment, including the coalescence formula and the coalescence determinant.
5.2. The pairwise coalescence problem
Given $n=2k$ particles at positions $x_{1}<\cdots<x_{n}$, we ask: what is the total weight of the pairwise coalescence event—the event that by time $T$, each consecutive pair $\{x_{1},x_{2}\}$, $\{x_{3},x_{4}\}$, $\ldots$, $\{x_{n-1},x_{n}\}$ has coalesced, meaning that the particles within each pair have merged (directly or through a chain of intermediate collisions) into a single heir? (Different pairs may merge separately or may share collisions.)
Computing this from a coalescence formula would require summing over all coalescence scenarios consistent with the required pairings—different coalescence trees and ghost configurations. We avoid this difficulty by converting the problem to annihilation.
5.3. The coalescence-to-annihilation map
The key observation is that pairwise coalescence can be reinterpreted as complete annihilation via a parity reclassification.
5.3.1. Parity of an entity
Define the parity of an entity as the parity of its multiplicity (Definition˜5.1): odd (representing an odd number of original particles) or even (representing an even number, including zero for ghosts).
5.3.2. Characterizing pairwise coalescence
The event “all $k$ consecutive pairs have coalesced” is equivalent to “every final heir has even multiplicity.” For the forward direction: each required pair contributes two original particles to the heir that absorbed it; if all pairs are absorbed, every heir’s multiplicity is even. For the converse: if every final heir has even multiplicity, then every original particle has merged with at least one other; since particles can only merge with neighbors and the consecutive pairing accounts for all $n=2k$ particles, each pair must have coalesced.
5.3.3. The reclassification
Given a coalescence performance where every final heir has even multiplicity, reclassify each entity as follows:
  • every entity with even multiplicity becomes a ghost;
  • every entity with odd multiplicity remains a particle.
At each binary coalescence step, the three cases are:
Incoming parities Outgoing parities Reclassified as
odd + odd even + zero ghost + ghost (annihilation)
odd + even odd + zero particle + ghost (non-event)
even + even even + zero ghost + ghost (non-event)
In the first case, two odd-parity entities (particles in the annihilation picture) produce two ghosts: this is an annihilation event. In the other cases, at least one incoming entity has even parity (already a ghost); the meeting is a non-event. In the third case, both incoming entities already have even parity (both are ghosts in the reclassified picture), so their meeting is likewise a non-event: ghosts pass through each other without interacting.
Since every final heir has even multiplicity, every final heir is reclassified as a ghost. The result is complete annihilation: zero survivors and $n=2k$ ghosts.
5.4. The Pfaffian formula
We now apply the annihilation formula to derive the Pfaffian structure.
5.4.1. The antisymmetric matrix
For particles $I<J$ (with initial positions $x_{I}\prec x_{J}$), define the pairwise annihilation weight
(5.1) $A_{IJ}=2\sum_{\begin{subarray}{c}a,b\\ a\prec b\end{subarray}}W(x_{I}\to b)\cdot W(x_{J}\to a)\;+\;\sum_{c}W(x_{I}\to c)\cdot W(x_{J}\to c),$
where both sums are over final vertices (replaced by integrals for continuous state spaces, in which case the second sum vanishes). The first sum is twice the total weight of strictly crossing paths: particle $I$ (starting left) ends strictly right of particle $J$ (starting right). The second sum is a diagonal correction for ghost pairs that end at the same vertex (for continuous state spaces such as Brownian motion, coincident endpoints have measure zero and this sum vanishes).
Extend to an $n\times n$ antisymmetric matrix by setting $A_{JI}=-A_{IJ}$ for $I<J$ and $A_{II}=0$.
Theorem 5.3 (Pfaffian pairwise coalescence formula).
For $n=2k$ particles at positions $x_{1}<\cdots<x_{n}$, the total weight of the pairwise coalescence event is
(5.2) $P=\operatorname{Pf}(A).$
Proof.
By the coalescence-to-annihilation map (Section˜5.3), the pairwise coalescence weight equals the total weight of complete annihilation of $n=2k$ particles. Apply the annihilation formula (Section˜3.3) with zero survivors and $k$ ghost pairs.
For any ghost sign pattern $(\varepsilon_{1},\ldots,\varepsilon_{k})$, expand $\det(M)$ via the Leibniz formula. With zero survivors, the bijection $\pi\colon\mathcal{A}\to\mathcal{G}$ assigns each actor to one of $k$ ghost pairs (two slots each); the Leibniz terms therefore group naturally by which actors share a ghost pair. The surviving terms are indexed by perfect matchings of $\{1,\ldots,n\}$.
Sum the annihilation formula over all ghost positions $(a_{j},b_{j})$ and all $2^{k}$ ghost sign patterns. When marginalizing over ghost positions, the ordering of ghost pairs becomes irrelevant: the physical outcome is an unordered collection of ghost pairs, and the marginalization treats all orderings equivalently. The result factorizes across pairs. For particles $I<J$ forming a ghost pair, the contribution summing both ghost signs is
$C_{IJ}=2\sum_{a\prec b}W(x_{I}\to b)\,W(x_{J}\to a)+\sum_{c}W(x_{I}\to c)\,W(x_{J}\to c)=A_{IJ}.$
Each perfect matching $\mu=\{\{p_{1},q_{1}\},\ldots,\{p_{k},q_{k}\}\}$ (unordered pairs) contributes $\prod_{l}C_{p_{l},q_{l}}$ with the standard Pfaffian sign (the sign of the permutation $(p_{1},q_{1},\ldots,p_{k},q_{k})$ where $p_{l}<q_{l}$ and $p_{1}<\cdots<p_{k}$).
Summing over matchings gives the Pfaffian:
$P=\operatorname{Pf}(A).\qed$
Example 5.4 (Four particles, two pairs).
Four particles at $x_{1}<x_{2}<x_{3}<x_{4}$ undergo pairwise coalescence: pair $\{x_{1},x_{2}\}$ and pair $\{x_{3},x_{4}\}$ each merge by time $T$. By parity reclassification, this is equivalent to complete annihilation of all four particles. The $4\times 4$ antisymmetric matrix $A$ has entries $A_{IJ}$ given by (5.1), and the Pfaffian expands as
$P=\operatorname{Pf}(A)=A_{12}\,A_{34}-A_{13}\,A_{24}+A_{14}\,A_{23}.$
The three terms correspond to the three perfect matchings of $\{1,2,3,4\}$. The matching $\{(1,3),(2,4)\}$ (sign 1-1) is not physically realizable: particles $x_{1}$ and $x_{3}$ cannot meet without $x_{2}$ intervening first (the consecutive collision property). The remaining matchings $\{(1,2),(3,4)\}$ and $\{(1,4),(2,3)\}$ (both sign +1+1) are realizable—the first when adjacent pairs annihilate directly, the second when $\{x_{2},x_{3}\}$ annihilate first and then $\{x_{1},x_{4}\}$ meet. However, each product $A_{ij}\,A_{kl}$ counts all path configurations where the two pairs meet, including ones with wrong collision order or unexpected crossings. The counterterm $A_{13}\,A_{24}$ cancels exactly these spurious contributions, by the sign-reversing involution.
5.5. Discussion
5.5.1. Annihilation proves coalescence
The proof uses the annihilation formula to establish a result about coalescence. This interplay between the two models—a payoff of the ghost framework—is made possible by the parity reclassification.
5.5.2. From determinant to Pfaffian
The annihilation formula is a single determinant. For complete annihilation, all columns are ghost-pair columns, and marginalizing over ghost positions produces a Pfaffian. With survivors, the formula remains a determinant (survivor columns contribute plain transition weights). The ghost structure thus interpolates between determinants (with survivors) and Pfaffians (without).
5.5.3. Connection to gap statistics
The Pfaffian pairwise coalescence formula connects to empty-interval probabilities and voter model duality. In the companion paper [Śni26], this connection is developed further: voter model duality bridges the ghost framework to Pfaffian point process theory [TZ11, GPTZ18], and the Pfaffian formula yields empty-interval probabilities for coalescing random walks.
Appendix A Prescribed annihilation: computational evidence
The annihilation formula (Section˜3.3) introduces ghost particles to restore the $n\times n$ matrix after annihilation has reduced the number of surviving particles, and labels the resulting ghost pairs by a uniform random numbering: they do not remember which initial particles produced them. Can the formula be refined to specify which particles annihilate—that is, can one obtain a Karlin–McGregor-type expression for the weight of a prescribed annihilation pattern? For $n=2$ the question is vacuous (there is only one pair), so the first nontrivial case is $n\geq 3$ with a single prescribed annihilation.
A.1. Prescribed annihilation
Definition A.1 (Prescribed annihilation).
Given $n$ particles at $x_{1}\prec\cdots\prec x_{n}$, fix adjacent particles $x_{m}$ and $x_{m+1}$. The prescribed annihilation weight $Z_{\text{prescribed}}$ is the total weight of all evolutions in which $x_{m}$ and $x_{m+1}$ annihilate (producing a ghost pair at positions $a\prec b$) while the remaining n2n-2 particles reach positions $y_{1}\prec\cdots\prec y_{n-2}$ satisfying $a\prec y_{1}\prec b$, without colliding with each other or with the annihilating pair before annihilation.
Since the ghost pair’s origin is determined by the constraint, the question is whether $Z_{\text{prescribed}}$ can be expressed as a linear combination of Karlin–McGregor products:
(A.1) $Z_{\text{prescribed}}=\sum_{\pi\in S_{n}}c_{\pi}\prod_{i=1}^{n}W(x_{i}\to z_{\pi(i)}),$
with rational coefficients $c_{\pi}$ independent of the positions and of time. Here $z_{1},\ldots,z_{n}$ denote the $n$ final positions (survivors and ghosts combined).
A.2. Computation
On the $\mathbb{Z}$ lattice (simple random walk with $\pm 1$ steps), the prescribed annihilation weight can be computed exactly by enumerating all $2^{nt}$ path evolutions and filtering to those satisfying the constraints of Definition˜A.1. For a fixed ordering of the $n$ final positions, distinct final-position tuples provide linear equations in the $n!$ unknowns $c_{\pi}$ of (A.1). The resulting system is solved over $\mathbb{Q}$ in exact arithmetic. (The same technique was used to discover the annihilation formula of the present paper and the coalescence formula of the companion paper [Śni26a]; it succeeds when a formula of the form (A.1) exists.)
For $n=3$ (so $n!=6$ unknowns), with initial positions $x_{i}=2(i-1)$ and annihilating pair $(1,2)$, at time $t=4$ the ordering $a\prec y_{1}\prec b$ yields $4$ distinct tuples $(a,y_{1},b)$:
$(-2,0,2),\quad(-2,0,4),\quad(-2,2,4),\quad(0,2,4).$
The $4\times 6$ system over $\mathbb{Q}$ is inconsistent: no rational coefficients $c_{\pi}$ satisfy all equations. This set of four tuples is minimal; any three of them admit a solution.
A.3. Interpretation
The inconsistency is not a failure of a specific ansatz such as a determinant or permanent: every expression of the form (A.1) is a linear combination of the $n!$ products $\prod_{i}W(x_{i}\to z_{\pi(i)})$, and the linear system tests all such combinations simultaneously. No rational coefficients $c_{\pi}$ exist. Restricting to the canonical ordering $a\prec y_{1}\prec b$ excludes formulas that might depend on the relative ordering of all final positions; the inconsistency persists.
Ghost anonymity is not merely a convenient bookkeeping device; the computation above suggests that it is essential for the existence of the formula.
Acknowledgments
We thank Maciej Dołęga, Sho Matsumoto, and Karol Życzkowski for stimulating discussions.
Claude Code (Anthropic) was used as an assistant during formula discovery and manuscript preparation.